OpenStudy (anonymous):

Mrs. West is 14 years younger than her aunt. If Mrs. West's age in years is as much below 60 as her aunt's age is over 40, how old is each?

OpenStudy (anonymous):

The above is not correct. Here is how to think of it: Aunt - Mr's West = 14 and Mr's West Aunt 60 - y 40 + y So, (40+y) - (60-y) = 14 40+y-60+y = 14 2y - 20 = 14 2y = 34 y = 17 This is not their age! Be careful Now, to find their age: Mr's West = 60-17=43 Aunt = 40+17 = 57 To check: 57-43 = 14 Here you go, solved =)

OpenStudy (anonymous):

I've deleted my incorrect answer.

OpenStudy (anonymous):

mlhuyck, is the answer incorrect? If you need more explanation, let me know.